OpenStudy (countrygirl1431):

The slope of the line containing the points (-5, 3) and (-2, 1) is -2/3 2/3 -3/2

OpenStudy (here_to_help15):

Slope-intercept form: y = mx + b, where m is the slope and b is the y-intercept Slope formula: m = (y2 - y1)/(x2 - x1), where (x2, y2) and (x1, y1) are points on the line Find the slope using the slope formula.

OpenStudy (here_to_help15):

Recognize the formula?

OpenStudy (countrygirl1431):

yes

OpenStudy (here_to_help15):

Think your big enough to handle this :)

OpenStudy (here_to_help15):

You can eliminate B most likely

OpenStudy (countrygirl1431):

i figured that much i put it down to A and C

OpenStudy (here_to_help15):

x - x1/x2 - x1 = y - y1 / y2 - y1.

OpenStudy (countrygirl1431):

thats where i get confused

OpenStudy (here_to_help15):

Ok so let me plug it in for you

OpenStudy (countrygirl1431):

thank you your an angle lol

OpenStudy (here_to_help15):

(-5, 3) and (-2, 1) \[x- \frac{- 5 }{-2-5 } = y - \frac{ 3 }{ 1-3 }\]

OpenStudy (here_to_help15):

angel?? lol

OpenStudy (here_to_help15):

OK so from there we just simply break it down let me do it for ya :) SO you can get a better understanding

OpenStudy (here_to_help15):

And oops the x and y is suppose to be in the fraction 00.00 at the top ehh as long as you undersand :)

OpenStudy (countrygirl1431):

yeah and your an angle is just something ive gotten into the habit of saying

OpenStudy (here_to_help15):

\[\frac{ x-5 }{ -7 } = \frac{ y-3 }{ -2 }\]

OpenStudy (here_to_help15):

I like it lol "angle" :)

OpenStudy (here_to_help15):

Now we cross multiply

OpenStudy (here_to_help15):

-2(x-5)=-7(y-3)

OpenStudy (here_to_help15):

Think you can handle or you still need help :)

OpenStudy (countrygirl1431):

hang on

OpenStudy (here_to_help15):

Ok

OpenStudy (countrygirl1431):

im still lost lol

OpenStudy (here_to_help15):

\[-2x+10= -7y+21\]

OpenStudy (here_to_help15):

NOW CAN YOU HANDLE IT!>!> lol

OpenStudy (countrygirl1431):

oh yeah

OpenStudy (countrygirl1431):

u dont have to yell geez

OpenStudy (here_to_help15):

lol

OpenStudy (here_to_help15):

\[-2x-11+7y=0\]

OpenStudy (here_to_help15):

to find the slope, write the equation in slope -intercept form i.e. y = mx+c.

OpenStudy (here_to_help15):

\[7y= 2x+11\]

OpenStudy (here_to_help15):

I think it might be B lol

OpenStudy (countrygirl1431):

idk

OpenStudy (here_to_help15):

\[y=\frac{ 2x }{ 7} + \frac{ 11 }{ 7}\]

OpenStudy (countrygirl1431):

im saying b

OpenStudy (here_to_help15):

Crap i crapped up :( @amistre64 check my work please?

OpenStudy (here_to_help15):

@acxbox22 hey bud check my work please i am wrong somewhere

OpenStudy (here_to_help15):

@Loser66

OpenStudy (here_to_help15):

@sammixboo

OpenStudy (here_to_help15):

@satellite73

OpenStudy (here_to_help15):

@TheSmartOne

OpenStudy (countrygirl1431):

lol

OpenStudy (here_to_help15):

No one comes !! :(

OpenStudy (here_to_help15):

@Nnesha

OpenStudy (countrygirl1431):

thats why im laughing lol

OpenStudy (here_to_help15):

lol

OpenStudy (here_to_help15):

@campbell_st

OpenStudy (loser66):

hahaha... let see what happen. ha!! it's a long.........

OpenStudy (here_to_help15):

lol sorry

OpenStudy (countrygirl1431):

have funn going thru it hahahaha

OpenStudy (here_to_help15):

lol

OpenStudy (here_to_help15):

I think i did the wrong thing but i was right if the question was stated a different way lol

OpenStudy (loser66):

Question: what are we supposed to do? find the slope of the line pass through (-5,3) and (-2,1)?

OpenStudy (acxbox22):